Am aircraft flies 800km due east and then 600km due north. Determine the magnitude of it's displacement
1
Expert's answer
2021-09-10T12:18:11-0400

"\text {let }\vec{N}(0,1)- \text{unit vector north}"

"\text {let }\vec{E}(1,0)- \text{unit vector east}"

"\vec N\perp \vec E"

"\text{Am aircraft flies 800km due east and then 600km due north.}"

"\vec D= 800*\vec E+600*\vec N"

"\vec D= (800,0)+(0,600)=(800,600)"

"|\vec D|=\sqrt{600^2+800^2}=1000km"


"\text{Answer:}1000km"